SOLVE USING THE SQUARE ROOT PROPERTY:
-2(x-2)^2= -6
HERE I'M CONFUSED ON WHAT ALL TO SQUARE??? DO I SQUARE EVERTHING ON THE RIGHT OR WHAT PLEASE HELP AS SOON AS YOU CAN... TAHNKS

Answer by checkley77(12569)

-2(x-2)^2=-6
divide both sides by -2
(x-2)^2=-6/-2
(x-2)^2=3
now take the square root of both sides
x-2=sqrt3
x=sqrt3+2
x=1.73+2
x=3.73 answer.